# InstDrive: Instance-Aware 3D Gaussian Splatting for Driving Scenes

### v1

@@ 1 Introduction
Reconstructing dynamic driving scenes from dashcam videos has emerged as a crucial task in autonomous driving, enabling scene-level understanding, behavior prediction, and interactive planning. While recent advances in neural rendering and 3D scene reconstruction—particularly those based on 3D Gaussian Splatting—have shown impressive results in modeling open-world outdoor environments blocknerf ; nsg ; unisim ; hugs ; streetgs , they largely focus on holistic scene reconstruction without explicit instance-level representations. This limitation hinders downstream applications such as object-level editing, controllable simulation, and fine-grained semantic understanding.
@@
The first class learns continuous per-point or per-Gaussian features, followed by post clustering (e.g., cosine similarity thresholding) to assign instance labels lerf ; language ; langsplat . However, such methods require sensitive hyperparameter tuning, suffer from limited clustering precision, and often fail to produce globally consistent instance IDs. The second category utilizes pre-processing techniques such as SAM-based tracking and ID propagation deva to provide consistent instance IDs across frames gsgroup . These methods are highly dependent on tracking quality and often fail under multi-camera view changes or extended video sequences, which are common in autonomous driving scenarios. The third line of work explores codebook-based vector quantization, which maps continuous features to discrete instance identities through learned codebooks opengs . These methods eliminate the need for pre-/post-processing and hold promise for instance labeling. Nevertheless, designing a robust and efficient codebook remains a fundamental challenge.
@@
In the continuous feature learning stage, we enforce 2D–3D consistency constraints to guide the instance representation. In 2D space, we utilize SAM sam masks to supervise contrastive learning: features within the same instance are pulled closer, while features across different instances are pushed apart. Meanwhile, in 3D space, we partition the point field into voxels and apply a voxel-based spatial consistency loss to mitigate the feature ambiguity introduced by the volumetric nature of Gaussian rendering.
@@ 2.1 Neural Rendering for Scene Reconstruction
Neural rendering has emerged as a key paradigm for 3D scene reconstruction, with Neural Radiance Fields (NeRF) nerf pioneering the use of volumetric MLPs for photorealistic novel view synthesis. However, its slow training and rendering speed significantly hinder its applicability in downstream tasks. To address this, many acceleration techniques have been proposed, including hash encoding, coarse-to-fine sampling, and multiscale representations instantngp ; mipnerf ; zipnerf , which have greatly improved rendering efficiency and scalability. Despite these advancements, the implicit volumetric representation of NeRF remains difficult to interpret and edit, limiting its suitability for interactive applications such as autonomous driving simulation. In contrast, 3D Gaussian Splatting 3dgs ; 2dgs ; gspro adopts an explicit and compact representation that models the scene as a set of anisotropic Gaussians. This structure enables real-time rendering while maintaining competitive visual quality. Given the strong real-time requirements of driving scenes and the need for editable 3D representations, we adopt 3D Gaussian Splatting as the foundational representation in our framework.
@@ 2.2 Reconstruction of Driving Scenes
Neural rendering has become central to autonomous driving scene reconstruction, enabling photorealistic digital twins from real-world sensor data. Extensions of NeRF to driving scenarios address challenges such as dynamic objects, large-scale environments, and multi-sensor inputs. For instance, NSG nsg models multi-object dynamics via scene graphs for object-level manipulation; MARS mars builds a modular, instance-aware NeRF that separates dynamic actors from static backgrounds; and UniSim unisim constructs neural feature grids from driving logs to support multi-sensor simulation and controllable scene synthesis. To further improve real-time performance and editability, 3D Gaussian-based methods have gained traction. Street Gaussians streetgs use semantic Gaussian primitives with tracked poses for real-time rendering and composable street scenes. HUGS hugs and OmniRe omnire enhance scalability and interactivity in urban reconstructions through advanced 3DGS techniques. However, most existing approaches still rely on unified scene representations or semantic segmentation, lacking explicit 3D instance-level decomposition. This limits fine-grained understanding and flexible manipulation of individual elements—capabilities that are essential for downstream tasks like autonomous driving, simulation, and digital twin construction.
@@
Recent studies on point cloud understanding have explored lifting 2D visual and linguistic priors into 3D via open-vocabulary learning. By aligning 2D foundation models (e.g., CLIP clip , DINO dino ) with point cloud features through contrastive training or multi-view fusion, these methods enable zero-shot 3D recognition and segmentation without 3D annotations ulip ; clip2point ; pointclipv2 ; lu2023open .
NeRF have been extended beyond photorealistic rendering to enable semantic scene understanding by incorporating semantics into implicit 3D representations. Several approaches adapt 2D segmentation models such as SAM sam into 3D by leveraging density-guided projection and cross-view fusion, enabling promptable segmentation and hierarchical grouping in 3D space sa3d ; san ; garfield . The explicit representation of 3DGS enables semantic extensions langsplat ; language ; featuregs , but most methods struggle with consistent instance ID mapping across views. Gaussian Grouping gsgroup and CoSSegGaussians cosseggaussians achieve global consistency through pre-processing, yet fail to generalize to long, multi-view driving sequences. OmniSeg3D omniseg3d and OpenGaussian opengs use contrastive learning without tracking data, but their complex pipelines limit scalability in open driving scenarios.
@@ 3.1 Preliminaries
3D Gaussian Splatting (3DGS) 3dgs provides a compact, explicit, and differentiable representation of 3D scenes, making it particularly suitable for instance-level understanding tasks. By associating each Gaussian with localized geometry, appearance, and additional learnable attributes, 3DGS naturally supports structured reasoning at the object level. This per-primitive expressiveness enables fine-grained semantic modeling and facilitates integration with point-based or region-based supervision signals. Gaussian Splatting represents a 3D scene as a set of $N$ anisotropic 3D Gaussians $\mathcal{G}=\{g_{i}\}_{i=1}^{N}$, where each Gaussian is parameterized as $g_{i}=(p_{i},s_{i},q_{i},\alpha_{i},c_{i})$, with 3D position $p_{i}\in\mathbb{R}^{3}$, scale $s_{i}\in\mathbb{R}^{3}$, orientation represented by a unit quaternion $q_{i}\in\mathbb{R}^{4}$, opacity $\alpha_{i}\in[0,1]$, and color $c_{i}$ modeled using spherical harmonics (SH) coefficients. For each pixel $u$, the set of overlapping Gaussians $\mathcal{N}_{u}$ is computed and sorted in front-to-back order. The final color $\mathbf{c}_{u}$ at pixel $u$ is composited using alpha blending:
@@ 3.2 Continuous Instance Feature Learning
When rendering the scene from a given viewpoint, each 3D Gaussian feature is projected into a 2D feature map, alongside the RGB image, such that each pixel $u$ obtains a feature vector $\mathbf{f}_{u}$. These 2D features are used to supervise the 3D representation via differentiable rendering. Since we do not perform cross-frame instance tracking and rely solely on per-frame masks without category labels, we lack ground-truth instance IDs across views. We adopt the contrastive loss formulation introduced in prior work omniseg3d ; contrastive ; opengs , leveraging its established effectiveness in promoting instance-aware feature clustering. The core insight is that pixels rendered from 3D Gaussians and falling within the same mask region should share the same object identity and thus have similar features, whereas features from different masks should be distinct to improve instance-level separability. Let $M^{1},M^{2},\dots,M^{N}$ denote the $N$ segmentation masks generated by SAM for a given training image. For each mask $M^{i}$, we compute its prototype feature $\bar{\mathbf{f}}^{i}$ by taking the average of the features of all pixels within that region. We then define a unified contrastive loss that combines both intra-mask consistency and inter-mask discrimination:
@@ 3.3 Quantized Instance Feature Learning
The learned feature representations are typically continuous, which are effective for general tasks but ill-suited for discrete instance-level decisions. Without ground-truth instance IDs, the challenge lies in partitioning these features into distinct, instance-aware clusters. A common approach is to compare query features with all instance features using metrics like cosine similarity, followed by threshold-based filtering. However, this method cannot produce global instance segmentation and is highly sensitive to threshold selection. Vector quantization offers an alternative solution. Early applications in 3DGS focus on compression, grouping similar Gaussians into shared codebook vectors to reduce memory and computation compgs ; lightgs . Recent work has explored its use for semantic understanding opengs , but challenges remain: codebook updates require costly clustering (e.g., k-means), and learned vectors often lack uniformity, leading to overlapping clusters, unstable training, and ambiguous instance assignments.
@@ 4.1 Settings
We conduct all experiments on full PandaSet pandaset , a large-scale multimodal dataset for autonomous driving that provides synchronized LiDAR and camera data with high-quality 3D annotations. It contains 103 scenes with diverse traffic conditions. For evaluation, we follow prior works opengs ; gsgroup and adopt mean Intersection over Union (mIoU) and mean Accuracy (mAcc) as the primary metrics across all classes. However, as both metrics are computed on projected 2D masks, higher scores do not necessarily indicate better 3D instance segmentation performance—thus, we also refer readers to our 3D qualitative results for a more comprehensive assessment.
@@
All experiments are conducted on a single NVIDIA RTX 4090 GPU. For a fair comparison, we follow the original GSGroup gsgroup setup by using DEVA deva to preprocess instance masks. In contrast, both our method and OpenGS opengs directly utilize masks generated by SAM without additional preprocessing. In our implementation, the feature dimension for each 3D Gaussian instance is set to 8. We train all scenes for 30,000 iterations, with the codebook module activated starting from the 10,000th iteration. To improve computational efficiency, the voxel-based consistency loss is computed every five iterations using a voxel grid size of 0.5 meters. We set the weights of the $\mathcal{L}_{\text{voxel}}$, $\mathcal{L}_{contra}$, $\mathcal{L}_{\text{pseudo}}$ to 0.1.

### v2

@@ 1 Introduction
Reconstructing dynamic driving scenes from dashcam videos has emerged as a crucial task in autonomous driving, enabling scene-level understanding, behavior prediction, and interactive planning. While recent advances in neural rendering and 3D scene reconstruction—particularly those based on 3D Gaussian Splatting—have shown impressive results in modeling open-world outdoor environments [1, 2, 3, 4, 5], they largely focus on holistic scene reconstruction without explicit instance-level representations. This limitation hinders downstream applications such as object-level editing, controllable simulation, and fine-grained semantic understanding.
@@
The first class learns continuous per-point or per-Gaussian features, followed by post clustering (e.g., cosine similarity thresholding) to assign instance labels [6, 7, 8]. However, such methods require sensitive hyperparameter tuning, suffer from limited clustering precision, and often fail to produce globally consistent instance IDs. The second category utilizes pre-processing techniques such as SAM-based tracking and ID propagation [9] to provide consistent instance IDs across frames [10]. These methods are highly dependent on tracking quality and often fail under multi-camera view changes or extended video sequences, which are common in autonomous driving scenarios. The third line of work explores codebook-based vector quantization, which maps continuous features to discrete instance identities through learned codebooks [11]. These methods eliminate the need for pre-/post-processing and hold promise for instance labeling. Nevertheless, designing a robust and efficient codebook remains a fundamental challenge.
@@
In the continuous feature learning stage, we enforce 2D–3D consistency constraints to guide the instance representation. In 2D space, we utilize SAM [12] masks to supervise contrastive learning: features within the same instance are pulled closer, while features across different instances are pushed apart. Meanwhile, in 3D space, we partition the point field into voxels and apply a voxel-based spatial consistency loss to mitigate the feature ambiguity introduced by the volumetric nature of Gaussian rendering.
@@ 2.1 Neural Rendering for Scene Reconstruction
Neural rendering has emerged as a key paradigm for 3D scene reconstruction, with Neural Radiance Fields (NeRF) [13] pioneering the use of volumetric MLPs for photorealistic novel view synthesis. However, its slow training and rendering speed significantly hinder its applicability in downstream tasks. To address this, many acceleration techniques have been proposed, including hash encoding, coarse-to-fine sampling, and multiscale representations [14, 15, 16], which have greatly improved rendering efficiency and scalability. Despite these advancements, the implicit volumetric representation of NeRF remains difficult to interpret and edit, limiting its suitability for interactive applications such as autonomous driving simulation. In contrast, 3D Gaussian Splatting [17, 18, 19] adopts an explicit and compact representation that models the scene as a set of anisotropic Gaussians. This structure enables real-time rendering while maintaining competitive visual quality. Given the strong real-time requirements of driving scenes and the need for editable 3D representations, we adopt 3D Gaussian Splatting as the foundational representation in our framework.
@@ 2.2 Reconstruction of Driving Scenes
Neural rendering has become central to autonomous driving scene reconstruction, enabling photorealistic digital twins from real-world sensor data. Extensions of NeRF to driving scenarios address challenges such as dynamic objects, large-scale environments, and multi-sensor inputs. For instance, NSG [2] models multi-object dynamics via scene graphs for object-level manipulation; MARS [20] builds a modular, instance-aware NeRF that separates dynamic actors from static backgrounds; and UniSim [3] constructs neural feature grids from driving logs to support multi-sensor simulation and controllable scene synthesis. To further improve real-time performance and editability, 3D Gaussian-based methods have gained traction. Street Gaussians [5] use semantic Gaussian primitives with tracked poses for real-time rendering and composable street scenes. HUGS [4] and OmniRe [21] enhance scalability and interactivity in urban reconstructions through advanced 3DGS techniques. However, most existing approaches still rely on unified scene representations or semantic segmentation, lacking explicit 3D instance-level decomposition. This limits fine-grained understanding and flexible manipulation of individual elements—capabilities that are essential for downstream tasks like autonomous driving, simulation, and digital twin construction.
@@
Recent studies on point cloud understanding have explored lifting 2D visual and linguistic priors into 3D via open-vocabulary learning. By aligning 2D foundation models (e.g., CLIP [22], DINO [23]) with point cloud features through contrastive training or multi-view fusion, these methods enable zero-shot 3D recognition and segmentation without 3D annotations [24, 25, 26, 27].
NeRF have been extended beyond photorealistic rendering to enable semantic scene understanding by incorporating semantics into implicit 3D representations. Several approaches adapt 2D segmentation models such as SAM [12] into 3D by leveraging density-guided projection and cross-view fusion, enabling promptable segmentation and hierarchical grouping in 3D space [28, 29, 30]. The explicit representation of 3DGS enables semantic extensions [8, 7, 31], but most methods struggle with consistent instance ID mapping across views. Gaussian Grouping [10] and CoSSegGaussians [32] achieve global consistency through pre-processing, yet fail to generalize to long, multi-view driving sequences. OmniSeg3D [33] and OpenGaussian [11] use contrastive learning without tracking data, but their complex pipelines limit scalability in open driving scenarios.
@@ 3.1 Preliminaries
3D Gaussian Splatting (3DGS) [17]provides a compact, explicit, and differentiable representation of 3D scenes, making it particularly suitable for instance-level understanding tasks. By associating each Gaussian with localized geometry, appearance, and additional learnable attributes, 3DGS naturally supports structured reasoning at the object level. This per-primitive expressiveness enables fine-grained semantic modeling and facilitates integration with point-based or region-based supervision signals. Gaussian Splatting represents a 3D scene as a set of $N$ anisotropic 3D Gaussians $\mathcal{G}=\{g_{i}\}_{i=1}^{N}$, where each Gaussian is parameterized as $g_{i}=(p_{i},s_{i},q_{i},\alpha_{i},c_{i})$, with 3D position $p_{i}\in\mathbb{R}^{3}$, scale $s_{i}\in\mathbb{R}^{3}$, orientation represented by a unit quaternion $q_{i}\in\mathbb{R}^{4}$, opacity $\alpha_{i}\in[0,1]$, and color $c_{i}$ modeled using spherical harmonics (SH) coefficients. For each pixel $u$, the set of overlapping Gaussians $\mathcal{N}_{u}$ is computed and sorted in front-to-back order. The final color $\mathbf{c}_{u}$ at pixel $u$ is composited using alpha blending:
@@ 3.2 Continuous Instance Feature Learning
When rendering the scene from a given viewpoint, each 3D Gaussian feature is projected into a 2D feature map, alongside the RGB image, such that each pixel $u$ obtains a feature vector $\mathbf{f}_{u}$. These 2D features are used to supervise the 3D representation via differentiable rendering. Since we do not perform cross-frame instance tracking and rely solely on per-frame masks without category labels, we lack ground-truth instance IDs across views. We adopt the contrastive loss formulation introduced in prior work [33, 34, 11], leveraging its established effectiveness in promoting instance-aware feature clustering. The core insight is that pixels rendered from 3D Gaussians and falling within the same mask region should share the same object identity and thus have similar features, whereas features from different masks should be distinct to improve instance-level separability. Let $M^{1},M^{2},\dots,M^{N}$ denote the $N$ segmentation masks generated by SAM for a given training image. For each mask $M^{i}$, we compute its prototype feature $\bar{\mathbf{f}}^{i}$ by taking the average of the features of all pixels within that region. We then define a unified contrastive loss that combines both intra-mask consistency and inter-mask discrimination:
@@ 3.3 Quantized Instance Feature Learning
The learned feature representations are typically continuous, which are effective for general tasks but ill-suited for discrete instance-level decisions. Without ground-truth instance IDs, the challenge lies in partitioning these features into distinct, instance-aware clusters. A common approach is to compare query features with all instance features using metrics like cosine similarity, followed by threshold-based filtering. However, this method cannot produce global instance segmentation and is highly sensitive to threshold selection. Vector quantization offers an alternative solution. Early applications in 3DGS focus on compression, grouping similar Gaussians into shared codebook vectors to reduce memory and computation [35, 36]. Recent work has explored its use for semantic understanding [11], but challenges remain: codebook updates require costly clustering (e.g., k-means), and learned vectors often lack uniformity, leading to overlapping clusters, unstable training, and ambiguous instance assignments.
@@ 4.1 Settings
We conduct all experiments on full PandaSet [37], a large-scale multimodal dataset for autonomous driving that provides synchronized LiDAR and camera data with high-quality 3D annotations. It contains 103 scenes with diverse traffic conditions. For evaluation, we follow prior works [11, 10] and adopt mean Intersection over Union (mIoU) and mean Accuracy (mAcc) as the primary metrics across all classes. However, as both metrics are computed on projected 2D masks, higher scores do not necessarily indicate better 3D instance segmentation performance—thus, we also refer readers to our 3D qualitative results for a more comprehensive assessment.
@@
All experiments are conducted on a single NVIDIA RTX 4090 GPU. For a fair comparison, we follow the original GSGroup [10] setup by using DEVA [9] to preprocess instance masks. In contrast, both our method and OpenGS [11] directly utilize masks generated by SAM without additional preprocessing. In our implementation, the feature dimension for each 3D Gaussian instance is set to 8. We train all scenes for 30,000 iterations, with the codebook module activated starting from the 10,000th iteration. To improve computational efficiency, the voxel-based consistency loss is computed every five iterations using a voxel grid size of 0.5 meters. We set the weights of the $\mathcal{L}_{\text{voxel}}$, $\mathcal{L}_{contra}$, $\mathcal{L}_{\text{pseudo}}$ to 0.1.